OpenStudy (astrophysics):

How do you graph r^2=sin2theta

OpenStudy (astrophysics):

@ganeshie8 \[r^2 = \sin(2 \theta)\]

OpenStudy (astrophysics):

\[\large r = \pm \sqrt{\sin 2\theta}\]

OpenStudy (astrophysics):

Nvm I got it, looks like made the sin2 theta graph and did the same thing earlier and got it :P.